# Sina Baharlouei

### 06 Dec 2023 v1

@@ 2.1 A Convergent Stochastic Algorithm for fair ERM via $f$-Divergences
$\min_{{\bm{\theta}}}\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\mathbb{P}_{s}(s=k)\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}})f\Big{(}\frac{\mathbb{P}_{\hat{y}_{{\bm{\theta}}},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)}{\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}}=j)\mathbb{P}_{s}(s=k)}\Big{)}$ ($f$-FERM)
While the non-linearity of $f$-divergences in ($f$-FERM) empowers the underlying model to capture more complex dependencies between sensitive attributes and predictions compared to the linear measures (Zafar et al., 2017), the objective function can no longer be represented as a summation of functions over input data points. Consequently, one cannot directly apply the stochastic gradient descent method (or its variations, such as Adam) to the objective function in ($f$-FERM). In particular, directly evaluating the gradient of the objective function of ($f$-FERM) on a mini-batch of data leads to a statistically biased estimation of the entire objective’s gradient. Such statistical biases prevent the convergence of algorithms such as SGD (even with a strongly convex minimization landscape) (Ajalloeian and Stich, 2020; Chen and Luss, 2018), let aside the more complex objectives arising in modern-day neural networks.
@@
$\min_{{\bm{\theta}}}\>\max_{{\mathbf{A}}}\quad\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}{\mathbf{A}}_{jk}\mathbb{P}_{\hat{y},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)-f^{*}({\mathbf{A}}_{jk})\mathbb{P}_{\hat{y}}(\hat{y}_{{\bm{\theta}}}=j)\mathbb{P}_{s}(s=k)\Big{]}$ (3)
@@
Proof  The proof is standard and appears in Appendix D.
@@
$\begin{split}\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}}=j)&=\frac{1}{n}\sum_{i=1}^{n}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})=\mathbb{E}\Big{[}\underbrace{\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})}_{\hat{\mathbb{P}}_{\hat{y}_{{\bm{\theta}}}}(j;\>\mathcal{B})}\Big{]}\\ \mathbb{P}_{\hat{y}_{{\bm{\theta}}},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)&=\frac{1}{n}\sum_{i=1}^{n}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)=\mathbb{E}\Big{[}\underbrace{\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)}_{\hat{\mathbb{P}}_{\hat{y}_{{\bm{\theta}}},s}(j,k;\>\mathcal{B})}\Big{]}.\end{split}$ (4)
@@
$\min_{{\bm{\theta}}}\>\max_{{\mathbf{A}}}\quad\frac{1}{n}\sum_{i=1}^{n}\left[\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}{\mathbf{A}}_{jk}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}({\mathbf{A}}_{jk})\pi_{k}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\Big{]}\right]$ (5)
Therefore, evaluating the gradient of the objective function with respect to the optimization variables ${\bm{\theta}}$ and ${\mathbf{A}}$ over a random small batch of data points leads to an unbiased gradient estimator of the entire objective function.
@@
4:   ${\bm{\theta}}^{t}={\bm{\theta}}^{t-1}-\frac{\eta_{\bm{\theta}}}{b}\sum\;\nabla_{\bm{\theta}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),y)-\eta_{\bm{\theta}}\lambda\nabla_{\bm{\theta}}\Big{(}{\mathbf{A}}_{jk}^{t-1}\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}},{\mathbf{s}}}(j,k;\>\mathcal{B}_{t})-\pi_{k}f^{*}({\mathbf{A}}_{jk}^{t-1})\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}(j;\>\mathcal{B}_{t})\Big{)}$
5:   ${\mathbf{A}}_{jk}^{t}={\mathbf{A}}_{jk}^{t-1}+\eta_{\alpha}\;\nabla_{{\mathbf{A}}}\left({\mathbf{A}}_{jk}^{t-1}\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}},{\mathbf{s}}}(j,k;\>\mathcal{B}_{t})-\pi_{k}f^{*}({\mathbf{A}}_{jk}^{t-1})\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}(j;\>\mathcal{B}_{t})\right)$
@@
Proof  The formal statement and proof are relegated to Appendix F.
Theorem 2 applies to all $f$-divergences listed in Table 1 for all batch-sizes (even as small as the batch size of $1$). More sophisticated algorithms can be used to obtain $\mathcal{O}(\epsilon^{-6})$ iteration complexity Rafique et al. (1810); Zhang et al. (2022). However, such algorithms use nested loops and require more hyperparameter tunings. We provide an example of such an algorithm in Appendix G. If the f-divergence leads to a strongly concave function in $\mathbf{A}$ or satisfies Polyak-Łojasiewicz condition (e.g., for $\chi^{2}$ divergence), a faster rate of $\mathcal{O}(\epsilon^{-5})$ can be obtained for this algorithm (Appendix F). In addition, if larger batch size of $\mathcal{O}(\epsilon^{-2})$ is used, we can further improve this rate to $O(\epsilon^{-4})$ iteration complexity (see Appendix F). Finally, when full batch size is used, then double/triple-loop algorithms can lead to the iteration complexity bounds of $O(\epsilon^{-2})$ in the nonconvex-strongly concave setting and $O(\epsilon^{-3})$ in the general nonconvex-concave setting; see (Kong and Monteiro, 2021; Nouiehed et al., 2019; Ostrovskii et al., 2021b; Thekumparampil et al., 2019).
@@
$\chi^{2}$ $(t-1)^{2}$ $\pi_{k}[{\mathbf{A}}_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-({\mathbf{A}}_{jk}+\frac{{\mathbf{A}}_{jk}^{2}}{4})\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]$
Reverse KL $-\ln t$ $\pi_{k}[{\mathbf{A}}_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}+(1+\ln(-{\mathbf{A}}_{jk}))\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]$
Total Variational $\frac{1}{2}|t-1|$ $\pi_{k}{\mathbf{A}}_{jk}[\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]\mathbb{I}_{\{|{\mathbf{A}}_{jk}|<1/2\}}$
KL $t\ln t$ $\pi_{k}[{\mathbf{A}}_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-e^{{\mathbf{A}}_{jk}-1}\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
Jensen-Shannon $-(t+1)\ln(\frac{t+1}{2})+t\ln t$ $\pi_{k}[{\mathbf{A}}_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}+\ln(2-e^{{\mathbf{A}}_{jk}})\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
Squared Hellinger $(\sqrt{t}-1)^{2}$ $\pi_{k}[{\mathbf{A}}_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}|{\mathbf{s}}_{k}}+({\mathbf{A}}_{jk}^{-1}+2)\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
@@ 3 Robust $f$-FERM in the Presence of Distribution Shifts
$\min_{{\bm{\theta}}}\>\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})\quad{\rm s.t.}\quad\max_{\mathbb{P}\in\mathcal{B}}\mathcal{D}_{f}\Big{(}\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s})||\>\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)\Big{)}\leq\delta.$ (6)
@@ 3.1 Robust $f$-FERM Under $\ell_{p}$ Norms and Small Distribution Shifts
$\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})\quad{\rm s.t.}\quad\max_{\begin{subarray}{c}||\mathbb{P}-\hat{\mathbb{P}}||_{p}\leq\delta\\ \begin{subarray}{c}||\mathbb{Q}-\hat{\mathbb{Q}}||_{p}\leq\delta\end{subarray}\end{subarray}}\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})\leq\kappa,$ (7)
@@
$\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\max_{\begin{subarray}{c}\|\mathbb{P}-\hat{\mathbb{P}}\|_{p}\leq\delta\\ \begin{subarray}{c}\|\mathbb{Q}-\hat{\mathbb{Q}}\|_{p}\leq\delta\end{subarray}\end{subarray}}\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})$ (8)
@@
$\min_{{\bm{\theta}}}\max_{\begin{subarray}{c}\|\mathbb{U}\|_{p}\leq\delta\\ \begin{subarray}{c}\|\mathbb{V}\|_{p}\leq\delta\end{subarray}\end{subarray}}\Bigg{(}h({\bm{\theta}},\mathbb{U},\mathbb{V}):=\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\langle\mathbb{U},\nabla_{\mathbb{P}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\rangle+\lambda\langle\mathbb{V},\nabla_{\mathbb{Q}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\rangle\Bigg{)},$ (9)
@@
$\min_{{\bm{\theta}}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\delta\|\nabla_{\mathbb{P}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\|_{q}+\lambda\delta\|\nabla_{\mathbb{Q}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\|_{q},$ (10)
where $\|\cdot\|_{q}$ is the dual of the $\ell_{p}$ norm with $\frac{1}{p}+\frac{1}{q}=1$. It can be shown that finding stationary points of (10) leads to a stationary point of (7):
Proposition 3
@@
A semi-stochastic memory-efficient first-order training algorithm. To apply (stochastic) gradient descent-ascent algorithm (Lin et al., 2020) to problem (9), we need to have unbiased estimator of the function $h({\bm{\theta}},\mathbb{U},\mathbb{V})$ w.r.t. ${\bm{\theta}}$, $\mathbb{U}$, and $\mathbb{V}$ variables. While it seems challenging to obtain unbiased estimator w.r.t. all variables, one can notice that if $\hat{p}_{j}({\bm{\theta}})$ and $\hat{q}_{j}({\bm{\theta}})$ can be computed easily with one forward pass over all data points (i.e., in $O(m\times n)$ memory requirement). Consequently, the gradient of $h({\bm{\theta}},\mathbb{U},\mathbb{V})$ w.r.t. $\mathbb{U}$ and $\mathbb{V}$ can be computed with one forward pass over all data points (without the need for doing backpropagation). On the other hand, one can easily obtain unbiased estimator of $\nabla_{{\bm{\theta}}}\hat{p}_{j}({\bm{\theta}})$ and $\nabla_{{\bm{\theta}}}\hat{q}_{j}({\bm{\theta}})$ in (11) using a small mini-batch of data. Such a task requires $O(b\times d)$ memory with $d$ being the number of parameters (i.e., ${\bm{\theta}}\in\mathbb{R}^{d}$) and $b$ being the batch size. Combining this unbiased estimation with the computed values of $\hat{p}_{j}({\bm{\theta}})$ and $\hat{q}_{j}({\bm{\theta}})$ leads to an unbiased estimator of the objective of (9) w.r.t. ${\bm{\theta}}$ variable. To summarize, we need to do one forward propagation to obtain gradients w.r.t. $\mathbb{U}$ and $\mathbb{V}$, and we only do backpropagation for computing gradients w.r.t. ${\bm{\theta}}$ over the mini-batch of data. Such an algorithm requires $O(mn+bd)$ memory requirement; and thus can be used for training large models (with $d,n\gg b,m$). It is known that memory requirements are the major limiting factors in training large models such as LLMs (Malladi et al., 2023).

### 07 Apr 2024 v2

@@ 2.1 A Convergent Stochastic Algorithm for fair ERM via $f$-Divergences
$\vspace{-2mm}\min_{{\bm{\theta}}}\quad\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\mathbb{P}_{s}(s=k)\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}})f\Big{(}\frac{\mathbb{P}_{\hat{y}_{{\bm{\theta}}},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)}{\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}}=j)\mathbb{P}_{s}(s=k)}\Big{)}$ ($f$-FERM)
While the non-linearity of $f$-divergences in ($f$-FERM) empowers the underlying model to capture more complex dependencies between sensitive attributes and predictions compared to the linear measures (Zafar et al., 2017), the objective function can no longer be represented as a summation of functions over input data points. Consequently, one cannot directly apply the stochastic gradient descent method (or its variations, such as Adam) to the objective function in ($f$-FERM). In particular, directly evaluating the gradient of the objective function of ($f$-FERM) on a mini-batch of data leads to a statistically biased estimation of the entire objective’s gradient. Such statistical biases prevent the convergence of algorithms such as SGD (even with a strongly convex minimization landscape) (Ajalloeian & Stich, 2020; Chen & Luss, 2018), let aside the more complex objectives arising in modern-day neural networks.
@@
$\vspace{-2mm}\min_{{\bm{\theta}}}\>\max_{A}\quad\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}A_{jk}\mathbb{P}_{\hat{y},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)-f^{*}(A_{jk})\mathbb{P}_{\hat{y}}(\hat{y}_{{\bm{\theta}}}=j)\mathbb{P}_{s}(s=k)\Big{]}$ (3)
@@
Proof.
The proof is standard and appears in Appendix D. ∎
@@
$\begin{split}\mathbb{P}_{\hat{y}_{{\bm{\theta}}}}(\hat{y}_{{\bm{\theta}}}=j)&=\frac{1}{n}\sum_{i=1}^{n}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})=\mathbb{E}\Big{[}\underbrace{\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})}_{\hat{\mathbb{P}}_{\hat{y}_{{\bm{\theta}}}}(j;\>\mathcal{B})}\Big{]}\\ \vspace{-1mm}\mathbb{P}_{\hat{y}_{{\bm{\theta}}},s}(\hat{y}_{{\bm{\theta}}}=j,s=k)&=\frac{1}{n}\sum_{i=1}^{n}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)=\mathbb{E}\Big{[}\underbrace{\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)}_{\hat{\mathbb{P}}_{\hat{y}_{{\bm{\theta}}},s}(j,k;\>\mathcal{B})}\Big{]}.\end{split}$ (4)
@@
$\min_{{\bm{\theta}}}\>\max_{A}\quad\frac{1}{n}\sum_{i=1}^{n}\left[\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\sum_{j\in\mathcal{Y},\atop k\in\mathcal{S}}\Big{[}A_{jk}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\mathbbm{1}(s_{i}=k)-f^{*}(A_{jk})\pi_{k}F_{j}({\mathbf{x}}_{i};{\bm{\theta}})\Big{]}\right]$ (5)
Thus, evaluating the gradient of the objective function w.r.t. the variables ${\bm{\theta}}$ and ${\mathbf{A}}$ over a random batch of data points leads to an unbiased estimator of the gradient of the objective function.
@@
4:   ${\bm{\theta}}^{t}={\bm{\theta}}^{t-1}-\frac{\eta_{\bm{\theta}}}{b}\sum\;\nabla_{\bm{\theta}}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),y)-\eta_{\bm{\theta}}\lambda\nabla_{\bm{\theta}}\Big{(}A_{jk}^{t-1}\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}},{\mathbf{s}}}(j,k;\>\mathcal{B}_{t})-\pi_{k}f^{*}(A_{jk}^{t-1})\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}(j;\>\mathcal{B}_{t})\Big{)}$
5:   $A_{jk}^{t}=A_{jk}^{t-1}+\eta_{\alpha}\;\nabla_{{\mathbf{A}}}\left(A_{jk}^{t-1}\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}},{\mathbf{s}}}(j,k;\>\mathcal{B}_{t})-\pi_{k}f^{*}(A_{jk}^{t-1})\hat{\mathbb{P}}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}(j;\>\mathcal{B}_{t})\right)$
@@
Proof.
The formal statement and proof are relegated to Appendix F. ∎
Theorem 2.2 applies to all $f$-divergences listed in Table 1 for all batch-sizes (even as small as the batch size of $1$). More sophisticated algorithms can be used to obtain $\mathcal{O}(\epsilon^{-6})$ iteration complexity Rafique et al. (1810); Zhang et al. (2022). However, such algorithms use nested loops and require more hyperparameter tunings. We provide an example of such an algorithm in Appendix G. If the f-divergence leads to a strongly concave function in $\mathbf{A}$ or satisfies Polyak-Łojasiewicz condition (e.g., for $\chi^{2}$ divergence), a faster rate of $\mathcal{O}(\epsilon^{-5})$ can be obtained for this algorithm (Appendix F). In addition, if larger batch size of $\mathcal{O}(\epsilon^{-2})$ is used, we can further improve this rate to $O(\epsilon^{-4})$ iteration complexity (see Appendix F). Finally, when full batch size is used, then double/triple-loop algorithms can lead to the iteration complexity bounds of $O(\epsilon^{-2})$ in the nonconvex-strongly concave setting and $O(\epsilon^{-3})$ in the general nonconvex-concave setting; see (Kong & Monteiro, 2021; Nouiehed et al., 2019; Ostrovskii et al., 2021b; Thekumparampil et al., 2019).
@@
$\chi^{2}$ $(t-1)^{2}$ $\pi_{k}[A_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-(A_{jk}+\frac{{\mathbf{A}}_{jk}^{2}}{4})\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]$
Reverse KL $-\ln t$ $\pi_{k}[A_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}+(1+\ln(-A_{jk}))\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]$
Total Variational $\frac{1}{2}|t-1|$ $\pi_{k}A_{jk}[\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}}]\mathbb{I}_{\{|A_{jk}|<1/2\}}$
KL $t\ln t$ $\pi_{k}[A_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}-e^{A_{jk}-1}\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
Jensen-Shannon $-(t+1)\ln(\frac{t+1}{2})+t\ln t$ $\pi_{k}[A_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{{\bm{\theta}}}|{\mathbf{s}}_{k}}+\ln(2-e^{A_{jk}})\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
Squared Hellinger $(\sqrt{t}-1)^{2}$ $\pi_{k}[A_{jk}\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}|{\mathbf{s}}_{k}}+(A_{jk}^{-1}+2)\mathbb{P}_{\hat{{\mathbf{y}}}_{\bm{\theta}}}]$
@@ 3 Robust $f$-FERM in the Presence of Distribution Shifts
$\vspace{-3mm}\min_{{\bm{\theta}}}\>\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})\quad{\rm s.t.}\quad\max_{\mathbb{P}\in\mathcal{B}}\mathcal{D}_{f}\Big{(}\mathbb{P}({\hat{y}_{{\bm{\theta}}}({\mathbf{x}}),s})||\>\mathbb{P}(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}))\otimes\mathbb{P}(s)\Big{)}\leq\kappa.$ (6)
@@ 3.1 Robust $f$-FERM Under $\ell_{p}$ Norms and Small Distribution Shifts
$\vspace{-1mm}\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})\quad{\rm s.t.}\quad\max_{\begin{subarray}{c}||\mathbb{P}-\hat{\mathbb{P}}||_{p}\leq\delta\\ \begin{subarray}{c}||\mathbb{Q}-\hat{\mathbb{Q}}||_{p}\leq\delta\end{subarray}\end{subarray}}\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})\leq\kappa,$ (7)
@@
$\vspace{-2mm}\min_{\bm{\theta}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\max_{\begin{subarray}{c}\|\mathbb{P}-\hat{\mathbb{P}}\|_{p}\leq\delta\\ \begin{subarray}{c}\|\mathbb{Q}-\hat{\mathbb{Q}}\|_{p}\leq\delta\end{subarray}\end{subarray}}\mathcal{D}_{f}(\mathbb{P}||\mathbb{Q})$ (8)
@@
$\vspace{-2mm}\min_{{\bm{\theta}}}\max_{\begin{subarray}{c}\|\mathbb{U}\|_{p}\leq\delta\\ \begin{subarray}{c}\|\mathbb{V}\|_{p}\leq\delta\end{subarray}\end{subarray}}\Bigg{(}h({\bm{\theta}},\mathbb{U},\mathbb{V}):=\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\langle\mathbb{U},\nabla_{\mathbb{P}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\rangle+\lambda\langle\mathbb{V},\nabla_{\mathbb{Q}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\rangle\Bigg{)},$ (9)
@@
$\vspace{-1mm}\min_{{\bm{\theta}}}\frac{1}{n}\sum_{i=1}^{n}\ell(\hat{y}_{{\bm{\theta}}}({\mathbf{x}}_{i}),y_{i})+\lambda\delta\|\nabla_{\mathbb{P}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\|_{q}+\lambda\delta\|\nabla_{\mathbb{Q}}\mathcal{D}_{f}(\hat{\mathbb{P}}||\hat{\mathbb{Q}})\|_{q},$ (10)
where $\|\cdot\|_{q}$ is the dual of the $\ell_{p}$ norm with $\frac{1}{p}+\frac{1}{q}=1$.
Proposition 3.1.
@@
A semi-stochastic memory-efficient first-order training algorithm. To apply (stochastic) gradient descent-ascent algorithm (Lin et al., 2020) to problem (9), we need to have unbiased estimator of the function $h({\bm{\theta}},\mathbb{U},\mathbb{V})$ w.r.t. ${\bm{\theta}}$, $\mathbb{U}$, and $\mathbb{V}$ variables. While it seems challenging to obtain unbiased estimator w.r.t. all variables, one can notice that if $\hat{p}_{j}({\bm{\theta}})$ and $\hat{q}_{j}({\bm{\theta}})$ can be computed easily with one forward pass over all data points (i.e., in $O(m\times n)$ memory requirement). Consequently, the gradient of $h({\bm{\theta}},\mathbb{U},\mathbb{V})$ w.r.t. $\mathbb{U}$ and $\mathbb{V}$ can be computed with one forward pass over all data points (without the need for doing backpropagation). On the other hand, one can easily obtain unbiased estimator of $\nabla_{{\bm{\theta}}}\hat{p}_{j}({\bm{\theta}})$ and $\nabla_{{\bm{\theta}}}\hat{q}_{j}({\bm{\theta}})$ in (11) using a small mini-batch of data. Such a task requires $O(b\times d)$ memory with $d$ being the number of parameters (i.e., ${\bm{\theta}}\in\mathbb{R}^{d}$) and $b$ being the batch size. Combining this unbiased estimation with the computed values of $\hat{p}_{j}({\bm{\theta}})$ and $\hat{q}_{j}({\bm{\theta}})$ leads to an unbiased estimator of the objective of (9) w.r.t. ${\bm{\theta}}$ variable. To summarize, we need to do one forward propagation to obtain gradients w.r.t. $\mathbb{U}$ and $\mathbb{V}$, and we only do backpropagation for computing gradients w.r.t. ${\bm{\theta}}$ over the mini-batch of data. Such an algorithm requires $O(mn+bd)$ memory requirement and thus can be used for training large models (with $d,n\gg b,m$). It is known that memory requirements are the major limiting factors in training large models such as LLMs (Malladi et al., 2023).